# Bounds and Bugs: The Limits of Symmetry Metrics to Detect Partisan Gerrymandering

### v1

@@ Abstract
We provide both a theoretical and empirical analysis of the Mean-Median Difference (MM) and Partisan Bias (PB), which are both symmetry metrics intended to detect gerrymandering. We consider vote-share, seat-share pairs $(V,S)$ for which one can construct election data having vote share $V$ and seat share $S$, and turnout is equal in each district. We calculate the range of values that MM and PB can achieve on that constructed election data. In the process, we find the range of vote-share, seat share pairs $(V,S)$ for which there is constructed election data with vote share $V$, seat share $S$, and $MM=0$, and see that the corresponding range for PB is the same set of $(V,S)$ pairs. We show how the set of such $(V,S)$ pairs allowing for $MM=0$ (and $PB=0$) changes when turnout in each district is allowed to be different.
Although the set of $(V,S)$ pairs for which there is election data with $MM=0$ is the same as the set of $(V,S)$ pairs for which there is election data with $PB=0$, the range of possible values for MM and PB on a fixed $(V,S)$ is different. Additionally, for a fixed constructed election outcome, the values of the Mean-Median Difference and Partisan Bias can theoretically be as large as 0.5. We show empirically that these two metric values can differ by as much as 0.33 in US congressional map data. We use both neutral ensemble analysis and the short-burst method to show that neither the Mean-Median Difference nor the Partisan Bias can reliably detect when a districting map has an extreme number of districts won by a particular party. Finally, we give additional empirical and logical arguments in an attempt to explain why other metrics are better at detecting when a districting map has an extreme number of districts won by a particular party.
@@ 1 Introduction
In this article, we study two partisan symmetry metrics that have been part of the academic study of gerrymandering for many decades, as well as on-the-ground expert testimony for many years: the Mean-Median Difference and the Partisan Bias. These metrics are embedded in publicly available websites to evaluate maps such as Dave’s Redistricting App [12] and PlanScore [15] , software such as Maptitude for Redistricting [5], and ensemble-creating python libraries such as Gerrychain [28]. While some excellent studies have revealed the “problems and paradoxes” of these kinds of metrics [10], they remain ubiquitous in software, libraries, and expert testimonies on partisan gerrymandering.
Here we conduct two different studies of these metrics. We first conduct a theoretical study in which, for fixed vote-share $V$ and seat share $S$, we calculate the possible range of values for the Mean-Median (MM) and Partisan Bias (PB) on constructed election data having vote share $V$ and seat share $S$ when turnout is equal in each district. As a corollary, we prove which vote-share, seat-share pairs $(V,S)$ have some constructed corresponding election outcome data with Mean-Median (MM) value 0. (A value of 0 for these metrics is a “perfect score,” indicating no gerrymandering). We do the same for Partisan Bias (PB), showing that, while the range of values for MM and PB are different for fixed $(V,S)$, the $(V,S)$ pairs allowing for $MM=0$ is the same as the $(V,S)$ pairs allowing for $PB=0$. We also show how those vote-share seat-share pairs change as voter turnout between districts changes.
Secondly, we conduct a series of empirical studies on the Mean-Median Difference and Partisan Bias. Although the set of $(V,S)$ pairs allowing for $MM=0$ is the same as the set of $(V,S)$ pairs allowing for $PB=0$, the values of the Mean-Median Difference and Partisan Bias can be different for a single election outcome. Indeed, the maximum theoretical difference between the two is 0.5, and a difference of up to 0.33 is seen in our empirical results, using current US districting maps.
We also empirically study the efficacy of these metrics through two techniques: neutral ensemble analysis, and a short burst analysis. Neutral ensemble analysis is a strategy used by many groups and researchers (see [20] and [21], for example) to detect gerrymandering. It involves creating a large number of valid redistricting maps (a neutral ensemble of maps, not chosen to optimize any value), and comparing a proposed map to the maps in the neutral ensemble. If the proposed map appears to be an outlier within the neutral ensemble, this may indicate the presence of gerrymandering. We show that both the Mean-Median Difference [^1] This is a replication study for the Mean-Median Difference, as it was first shown in [30]. The results for Partisan Bias, however, are new. and the Partisan Bias are not as effective at detecting maps with an extreme number of districts won by some party as other metrics when used on a neutral ensemble.
The Short Burst method was developed by Cannon et al [8] to create “extreme maps” (in their article, they focus on maps with extreme numbers of majority-minority districts). We use their method to produce maps with extreme numbers of Republican-won districts, and maps with extreme numbers of Democratic-won districts. We use the maps created through the Short Burst method to show that for both the Mean-Median Difference [^2] Again, this is a replication study for the Mean-Median Difference, as it was first shown in [30], and the results for Partisan Bias are new. and Partisan Bias, extreme values of those metrics do not correspond to maps with an extreme number of Republican-won districts (or an extreme number of Democratic-won districts).
1.1 Background and Literature Review
@@
Partisan gerrymandering has recently become a rich topic of study for mathematicians and computer scientists, and the most straightforward and basic tools that have been proposed to detect gerrymandering are metrics. Historically, shape metrics such as the Polsby-Popper, Reock, and Convex Hull measures have been used to measure the irregularity of a district’s shape (see, for example, [1]). But recent studies have shown that having reasonable shape metric scores does not preclude a map from being an extreme partisan gerrymander (see, for example, [22]), and that even calculating shape metric values is subject to manipulation and abuse [4].
In the 1970s and for decades after, measures of partisan symmetry such as the Mean-Median Difference and Partisan Bias arose in the academic and legal communities [29, 16, 13, 23, 18, 26]. Symmetry metrics hinge on the idea that a map that gives party $A$ seat share $S$ when the vote share is $V$ must also give party $B$ seat share $S$ (or approximately $S$) when party $B$’s vote share is $V$ in order to not be a gerrymander. Other metrics such as the GEO metric [6], Efficiency Gap [31], and Declination [36] have been proposed in more recent years.
Of course, each of these metrics has their own shortcomings. In [35], Warrington compared the Partisan Bias, Mean-Median Difference, Efficiency Gap, and Declination on 12 hypothetical election outcomes, arguing that his Declination had fewer false positives and false negatives on those 12 constructed examples. In [10], DeFord et al showed that partisan symmetry metrics such as the Mean-Median Difference and Partisan Bias can be unable to distinguish between Republican and Democratic favoring maps on a couple of sample states, and argue that, “For those who do want to constrain the most extreme partisan outcomes that line drawing can secure, these investigations should serve as a strong caution regarding the use of partisan symmetry metrics.” Veomett [34] showed that the Efficiency Gap does not satisfy the Efficiency Principle (as its creators claimed [32]). Campisi at all [7] showed that the Declination does not catch all “packing and cracking” (the techniques understood to implement a gerrymander) as its creator claimed [36]. And in [30], Ratliff et al compared the Partisan Bias, Efficiency Gap, Declination, and GEO metrics in terms of their gameability. They argued that no metric is meaningfully different from measuring “districts won by party $A$” when paired with a neutral ensemble analysis, and that the Mean-Median difference had particularly unreliable and misleading results.
In this article, we fill a significant knowledge gap for the Mean-Median Difference and Partisan Bias metrics. Specifically, given a metric $m$, a very natural question is: For which vote-share, seat share pairs $(V,S)$ do there exist election data with vote share $V$, seat share $S$, and metric $m$ indicates no gerrymandering? This question has been answered for the Efficiency Gap [34] and the Declination [7]. Here, we answer that previously unanswered question for the Mean-Median Difference and the Partisan Bias. Indeed, we go even further by showing the potential range of values for MM and PB for a fixed vote-share, seat-share pair $(V,S)$.
We will also expand the empirical studies on symmetry metrics mentioned above. The authors of [10] considered three states, with a single election outcome used in the partisan analysis for each state. The authors of [30] used 18 different maps (3 maps for each of 6 states), but a single election outcome for each state. Here, we use the same maps considered in [30] but consider three elections per state in our neutral ensemble analysis study. We also add the partisan bias, which was not considered in [30]. Through these additional studies, we affirm the conclusion in [10] quoted above (that the Mean-Median Difference and Partisan Bias cannot constrain extreme partisan outcomes), and show that Partisan Bias is as erratic as the Mean-Median Difference was shown to be in [30].
We add additional empirical studies showing how MM and PB can differ in real-world data. We also provide logical and empirical arguments for why the Efficiency Gap, Declination, and GEO metric track “districts won” better than MM and PB.
1.2 “Districts Won” as the Foundation of Gerrymandering
The ultimate goal of a gerrymandering mapmaker is focused on the districts that are won (or projected to be won) by some party or group. It is for this reason that we focus our neutral ensemble analysis and short burst analysis on a metric’s ability to indicate when a party is winning an extreme number of districts in a state.
It may be the case that a metric is able to detect an extreme number of districts won, and also detect something additional about a map. For example, we will see that the GEO metric is effective at detecting extreme maps, and others have argued that the GEO metric is able to give additional information about which districts or regions of the state are potentially gerrymandered [6].
It may also be the case that a metric is unable to detect an extreme number of districts won, but detects something else entirely. We argue that this is the case for the Mean-Median Difference and Partisan Bias. As described in Section 1.3, Section 1.4, and [10], these two metrics are symmetry metrics that can show the asymmetry in outcomes for the two parties under an assumption of uniform partisan swing. However, as we shall see, these metrics are not reliable at detecting an extreme number of districts won. Since the districts won is the focus of a gerrymandering mapmaker, we argue that this implies that MM and PB are not effective at detecting gerrymandering.
1.3 Definitions

### v3

@@ Abstract
We consider two symmetry metrics commonly used to analyze partisan gerrymandering: the Mean-Median Difference (MM) and Partisan Bias (PB). Our main results compare, for combinations of seats and votes achievable in districted elections, the number of districts won by each party to the extent of potential deviation from the ideal metric values, taking into account the political geography of the state. These comparisons are motivated by examples where the MM and PB have been used in efforts to detect when a districting plan awards extreme number of districts won by some party. These examples include expert testimony, public-facing apps, recommendations by experts to redistricting commissions, and public policy proposals.
To achieve this goal we perform both theoretical and empirical analyses of the MM and PB. In our theoretical analysis, we consider vote-share, seat-share pairs $(V,S)$ for which one can construct election data having vote share $V$ and seat share $S$, and turnout is equal in each district. We calculate the range of values that MM and PB can achieve on that constructed election data. In the process, we find the range of $(V,S)$ pairs that achieve $MM=0$, and see that the corresponding range for PB is the same set of $(V,S)$ pairs. We show how the set of such $(V,S)$ pairs allowing for $MM=0$ (and $PB=0$) changes when turnout in each district is allowed to vary. By observing the results of this theoretical analysis, we can show that the values taken on by these metrics do not necessarily attain more extreme values in plans with more extreme numbers of districts won. We also analyze specific example elections, showing how these metrics can return unintuitive results. We follow this with an empirical study, where we show that on 18 different U.S. maps these metrics can fail to detect extreme seats outcomes.
@@ 1 Introduction
Partisan gerrymandering has recently become a rich topic of study for mathematicians and computer scientists. The most straightforward and basic tools that have been proposed to detect gerrymandering are metrics that associate a single numerical value with a given combination of districting plan and election data. Historically, shape metrics such as the Polsby-Popper, Reock, and Convex Hull metrics have been used to measure the irregularity of a district’s shape (see, for example, [1]). But recent studies have shown that having reasonable shape metric scores does not preclude a map from being an extreme partisan gerrymander (see, for example, [33]), and that even calculating shape metric values is subject to manipulation and abuse [4].
Thus, the focus on metrics to detect gerrymandering has shifted towards metrics which do not depend on district shapes. There has been a flurry in the past decade to produce such metrics, which include the Efficiency Gap [46], Declination [51], and GEO metric [8]. The Mean-Median Difference and Partisan Bias are two of the older (and thus arguably more well-established) metrics [^1] All of these non-shape metrics just named (the Efficiency Gap, Declination, GEO metric, Partisan Bias, and Mean-Median Difference) are designed to be used for a two-party system. In our theoretical discussions, we will thus refer to “party $A$” and “party $B$.” In our empirical results we consider two-party vote shares for the Democratic party and the Republican party., falling under the family of partisan symmetry metrics. Symmetry metrics hinge on the idea that gerrymandering happens when the two parties are not treated equally under shifts in the underlying vote data. For example, suppose a map gives party $A$ seat share $S$ when the vote share for party $A$ is $V$. Symmetry metrics have values far from 0 (indicating gerrymandering) when this same map would award party $B$ a seat share very different from $S$ when party $B$’s vote share is $V$ [^2] The sign of the partisan symmetry metric is intended to indicate which party is being advantaged; presumably this sign should indicate that party $A$ is advantaged if party $B$ would win a seat share less than $S$ when party $B$’s vote share is $V$, and the sign should indicate that party $B$ is advantaged if party $B$ would win a seat share more than $S$ when party $B$’s vote share is $V$.. The desire for partisan symmetry was highlighted by Niemi and Deegan in [42]. Grofman, King, Gelman, and others [27, 24, 35, 29, 34] have studied and promoted the Partisan Bias metric; the Mean-Median Difference was explicitly discussed as a metric by McDonald and Best in [38] and by Nagle in [41]. See the extensive information presented in [15] for additional historical and practical discussion of these metrics.
In [15] these metrics were analyzed from a theoretical perspective in terms of the ‘gaps’ between the values in the sorted vote vector and related to the Partisan Gini score analyzed in [34]. The main theoretical result of [15] focused on characterizing the cases where where the Partisan Gini (and hence MM and PB) metric is zero and then gave a series of paradoxes and empirical examples showing that the results of these metrics can be unintuitive or uncorrelated with seats outcomes. This paper extends the analysis and results of [15] by considering the behavior of the metrics across their full range of potential values.
As evidenced above, the Mean-Median Difference and Partisan Bias have been part of the academic study of gerrymandering for many decades. They have also been used in on-the-ground expert testimony for many years. These metrics are embedded in publicly available websites to evaluate maps such as Dave’s Redistricting App [23] and PlanScore [26], and software such as Maptitude for Redistricting [7]. They can be calculated using built-in functions in Python libraries that are used by researchers to study gerrymandering such as GerryChain [40]. While some excellent studies have revealed the “problems and paradoxes” of these kinds of metrics [15], they remain ubiquitous as established metrics to detect gerrymandering.
1.1 Contributions
In this article, we describe the bounds of values that these metrics can take on, and the “bugs” that these metrics exhibit in detecting partisan gerrymandering as measured by the number of districts in which a party is favored. Specifically, the framework we apply is to evaluate a metric’s ability to suggest when a districting plan has an extreme number of districts won by a particular party. While this is often the way the general public understands gerrymandering, and is also the focus of many experts and researchers, we note that there are many other types of gerrymandering, such as incumbency protection or anti-competitive gerrymanders, that are not directly addressed by this analysis. However, we believe that understanding and modeling this relationship is important, and in Section 2.2, we point out many instances in which the Mean-Median Difference and Partisan Bias are used for that purpose, such as expert testimony, public-facing apps, recommendations to independent redistricting commissions, and public policy.
With this framework in place, we present two types of studies; a theoretical analysis of the values that can be taken on by these metrics for a fixed vote-share, seat-share pair $(V,S)$ (the “bounds” study), and an empirical study of the actual values that these metrics take on for potential redistricting maps, using a fixed election as partisan data (the “bugs” study). In our bounds study, we point out the ways in which the values taken on by these metrics run counter to the expectation that maps with more extreme numbers of districts won should be able to exhibit more extreme values on these metrics. And in our bugs study, we show that these metrics do not take on more extreme values in maps with more extreme numbers of districts won.
We then recommend that the Mean-Median Difference and Partisan Bias not be used as the sole metric to detect gerrymandering in settings such as public facing apps, by line drawing commissions, court expert reports, and in public policy discussions. In particular, we recommend that they not be used as a proxy for “extreme number of districts won.”
We provide the motivation for this analysis and the main assumptions we are operating under in Section 2. We then give the mathematical definitions of the Mean-Median Difference and Partisan Bias in Section 3, which also summarizes other previously known facts about these metrics. We give the mathematical results (the “bounds” study) in Section 4; the proofs of those results are in Appendix A. The empirical results (the “bugs study”) are in Section 5.
2 Motivation and Background
In this section we provide motivation for this work by describing situations in which the partisan symmetry measures are used in practice to evaluate and optimize redistricting plans. We also point out that the Mean-Median Difference and Partisan Bias are ubiquitously used in an effort to describe situations where a party wins a larger than expected number of districts. This provides some context for our work comparing the results of partisan symmetry metrics on a fixed districting plan to the number of seats won by a party under that plan. Specifically, throughout this paper we focus on the ability of a metric to reflect situations where an extreme number of districts is won by a particular party as motivated by the use cases described below. Much of the background material below is motivated by the thorough coverage in [15], which discusses many of the same issues.
2.1 Counting Districts Won
That the focus of a partisan gerrymander is to create an extreme number of districts won by one party is a point of agreement among quantitative social scientists, political scientists, mathematicians, and the general public. In their recent article [34], Katz, King, and Rosenblatt state that “Partisan gerrymanderers use their knowledge of voter preferences and their ability to draw favorable redistricting plans to maximize their party’s seat share.” They go on to emphasize that a gerrymanderer cares not about issues such as turnout or electoral responsiveness “unless it helps them win more seats.” In his “Defining a Gerrymander” section of [30], Grofman explains that gerrymandering results in one party winning an extreme number of seats compared to the number they might win in “a plan drawn on the basis of neutral principles.” In [49], Wang describes the effects of partisan gerrymandering on the US House in 2012 by the number of seats won by each party, and states that “In summary, the effects of partisan redistricting exceeded the amount of asymmetry caused by natural patterns of population.” And in [15], DeFord et all state that “the express intent of a partisan gerrymander is to secure for their party as many seats as possible under the constraints of voter geography and the other rules of redistricting.” While several of these authors caveat that what counts as “extreme” may depend on a region’s political geography [^3] We acknowledge that a state’s political geography impacts what specific numbers may be considered “extreme number of districts won.” This is why we conduct our empirical analysis in Section 5; in that section, we produce extreme maps using a state’s political geography., they all agree that number of districts won is the foundation of a partisan gerrymander.
The general public agrees. News stories on gerrymandering virtually always cite the extreme number of districts won as evidence of a gerrymander (see [22], for example). Of course, some who seek to draw maps to benefit their political party, and some of those who study such maps, may take other factors into consideration such as protecting incumbents or creating partisan firewalls [^4] A partisan firewall can be created by, for example, ensuring a durable majority in a legislative body without necessarily maximizing the number of seats won.. We agree that those strategies also disenfranchise voters and that a gerrymandering mapmaker may draw the map also aiming at those goals. However, incumbency protection and partisan firewalls are not the first things that the general public and and many researchers consider when looking to identify and analyze partisan gerrymandering. For the purposes of this article, we take the common position of the above scholars (and many others) that, in order to detect partisan gerrymandering, we must, first and foremost, be able to detect an extreme number of districts won.
We also agree with Wang [49] and DeFord [15] that the political geography of a region being redistricted has significant impact on what number of districts won should be considered “extreme.” For example, Duchin et al showed that it was virtually impossible for Massachusetts to have a US House District with a majority of Republican voters, due to the political geography of that state [19]. Thus, a result of 9 out of 9 seats for Democrats should not be flagged as extreme at all for Massachusetts. We take into account each state’s political geography in our empirical analyses in Section 5.
For the purposes of our analysis, we need partisan data in order to have an estimate, for a potential redistricting map, as to how many seats each party could win with that map. To do this, we use past election data as a proxy for partisan preference. Using a single election outcome gives a clear interpretation of the partisan data, as opposed to an index which combines a variety of different data. This is why, in our analyses, we refer to the number of districts “won:” we report the number of districts that would have been won with a potential map, using partisan data from a past statewide election. In this paper we address the question of when partisan symmetry metrics do or do not effectively represent the number of districts won by a party, and we are particularly interested in the situation where extreme values in the metrics do not correspond to extreme values in the number of seats won and vice versa.
2.2 Motivating Examples
A key reason for focusing on MM and PB in this article is that these metrics have found common usage among the community of experts analyzing redistricting plans. These use cases have become quite broad and have begun to be adopted by members of the public, including through public-facing apps and in commissions. In these settings, much of the nuance of the metrics can be lost, leading to individuals focusing on simply minimizing these values as a goal to reduce vote dilution or to increase fairness. We argue in Sections 4 and 5 that this approach is unlikely to be generically successful, but in this section we provide some examples of the ways these metrics are being used in practice to support the need for a closer analysis of their consequences. One area of concern here is that individuals will attempt to optimize for these metrics and claim that any map reaching the values close to zero must be fair. Given the range of seats outcomes possible with these metric scores, this is a potentially misleading conclusion.
2.2.1 Expert Testimony
Both the Partisan-Bias and Mean-Median difference have become ubiquitous in expert reports analyzing redistricting plans for court cases. This includes experts writing in amicus briefs regarding gerrymandering, including LULAC v. Perry (2006) [36], Whitford v. Gill (2018) [25], and Rucho v. Common Cause (2019) [28], as well as by testifying experts in their reports including Session v. Perry (2003) [2], Perez v Perry (2011) [39], Favors v. Cuomo (2012) [3], and an appeal of Clarke v. Wisconsin Elections Commission (2023) [14, 13]. In court cases where many parties have proposed maps to be evaluated, it is often the case that the experts representing each party compute and discuss the symmetry measures. As an example, in the Congressional and Legislative redistricting lawsuits in Pennsylvania in 2022, over half a dozen experts used these values in their analyses [12]. These examples and the judicial opinions relying on them highlight just how ubiquitous these measures have become for evaluating redistricting plans.
In addition to the amicus briefs and reports listed above, court consultants and special masters have also relied on these metrics to analyze redistricting plans. A significant example from 2024 includes the analysis of Cervas and Grofman as court appointed consultants in Wisconsin [11]. In their report they analyze both the MM and PB across a variety of elections to determine whether there is bias in the proposed maps. As an example, after measuring the PB for all of the plans across 13 elections they write:
The deviations from political neutrality in these maps are a significant reduction from the Current plan and are similar to values that other state courts have viewed as acceptable compliance with their state constitution regarding neither favoring nor disfavoring a particular party (though 0 bias is preferable).
This type of statement directly incentivizes line drawers to attempt to minimize PB, perhaps across a set of available election data, in order to construct maps achieving this ‘preferable’ value. Regardless of what the creators and promoters of the Mean-Median Difference and Partisan Bias metrics intended those metrics to detect and measure, a takeaway from these reports in the court of public opinion is that they detect gerrymandering and by proxy whether an unfair number of districts would be won by one party under the plan under consideration.
2.2.2 Public-Facing Apps
Several apps have recently become available that make it easier than ever for the public to engage with the line drawing process, including drawing their own maps and evaluating current and proposed maps in their states. Examples of a line drawing app include Dave’s Redistricting App [23], while PlanScore has become a popular tool for scoring enacted and proposed redistricting plans[26]. The scores are also available to assess maps in the Maptitude for Redistricting software [7]. These apps offer friendly interfaces and results, and each of them reports on the partisan symmetry measures, often describing them as effective ways to detect unfairness or gerrymandering, and complementarily as optimization targets for fair maps. For example, the Dave’s Redistricting App blog about Partisan Bias [44] writes that “Because larger values mean more bias, smaller raw values of partisan bias are better.” This perspective is often adopted by individuals that use these apps, although the app developers themselves note that there are significant complexities in comparing proportionality to partisan symmetry scores [43].
Additionally, the Princeton Gerrymandering Project [50] and the Harvard-based ALARM Project [37] have recently released ensemble analyses of most states in the country. Most of the “Redistricting Report Cards” that can be found on the Princeton Gerrymandering Project’s site [50] report both the Mean-Median Difference and Partisan Bias, and the ALARM Project reports the Partisan Bias and Efficiency Gap scores for its ensembles. The tooltip on the PGP report card for the Mean-Median Difference states that “If the difference is large it could indicate a gerrymandered map where voters are crammed in a few districts.” As with the apps, these comments lead the audience to the conclusion that smaller values lead to better representational outcomes.
2.2.3 Optimizing for Mean-Median
Another formal setting in which the partisan symmetry metrics were analyzed was during the line drawing commissions work in Michigan. In the report, “A Commissioners Guide to Redistricting in Michigan,” prepared by researchers at Princeton University [5], a clear directive is given: “Commissioners should make an effort to minimize the degree of lopsided wins, mean-median difference, and the efficiency gap while being consistent with other higher-ranked criteria.” In this case a low MM value was sought as an ideal property of a plan, although the analysis in [15] shows that this is not an effective approach on its own. The Mean-Median Difference was used in an expert report to the Michigan Independent Citizens Redistricting Commission [32], and news reports suggest that the Michigan Independent Citizens Redistricting Commission was recommended to use maps having the Mean-Median Difference within symmetric bounds around 0, with the expectation that this would limit partisan gerrymandering [21].
Additionally, partially fueled by the rise of the public-facing apps discussed above, this redistricting cycle saw a large surge in map submissions from the general public. Sometimes in response to map making competitions, these map makers often argued for the adoption of their plans based on scores computed from Dave’s Redistricting App. An example of this is the proposals and amicus briefs filed by Petering in the Wisconsin redistricting litigation discussed above. In these briefs [13] Petering argued for the adoption of exactly the metrics reported by Dave’s Redistricting App and used those metrics to advance arguments about the maps under consideration. The Wisconsin court case where Petering filed his maps was particularly interesting, as the parties were required to upload copies of their maps to Dave’s Redistricting App upon submission at the request of the court appointed consultants quoted above.
2.2.4 Public Policy
Recent legislative efforts in the United States, at both the state and federal levels, have attempted to incorporate standards based on partisan symmetry into criteria for evaluating plans and preventing gerrymandering. Consider the following text from H.R. 1 as passed by the House of Representatives in 2021:
   [...] a redistricting plan shall be deemed to have the effect of unduly
                favoring or disfavoring a political party if--
                            (i) modeling based on relevant historical
                        voting patterns shows that the plan is
                        statistically likely to result in a partisan
                        bias of more than one seat in States with 20 or
                        fewer congressional districts or a partisan
                        bias of more than 2 seats in States with more
                        than 20 congressional districts, as determined
                        using quantitative measures of partisan
                        fairness, which may include, but are not
                        limited to, the seats-to-votes curve for an
                        enacted plan, the efficiency gap, the
                        declination, partisan asymmetry, and the mean-
                        median difference, [...]
This very clearly situates the Partisan Bias and Mean-Median Difference as metrics for detecting gerrymandering and hence could incentivize line drawers to minimize these quantities in order to pass the relevant test. See further discussion in [20] of the relationship between the test proposed in the Freedom to Vote Act and partisan symmetry measures, including in terms of their normative properties.
Legislation addressing partisan fairness measures in redistricting has also been proposed at the state level, particularly in states with a citizen initiative process. There has often been a lack of modeling in advance of the adoption of these metrics, which can lead to unintended consequences when line drawers attempt to optimize for these ostensibly neutral measures [15, 17]. These conflicts are often unintuitive to the public and can lead to partisan tension focused on language. The public discussion around the proportionality requirement and corresponding failure of Ohio’s Issue 1 provides an example of this type of conflict.
2.2.5 Motivating Examples: Summary
In each of these domains the MM and PB, among other partisan symmetry metrics, are being used to support the analysis of districting plans and to detect or mitigate gerrymandering. In many of these cases, there are unstated assumptions about the relationship between idealized values of the metrics and the common public understanding of gerrymandering in terms of the number of seats won by each party. Throughout this paper and particularly in Section 4.3 we use these practical examples to motivate our results.
3 Definitions and other Established Facts